How to solve this equation 5y+x=40 y=3x+40 with substitution
Answer by TimothyLamb(4379)

---
5y + x = 40
y = 3x + 40
---
put the system of linear equations into standard form
---
x + 5y = 40
3x - y = -40
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x= -10
y= 10
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php